# A Scalable Optimization Approach for Equitable Facility Location: Methodology and Transportation Applications

### v3

@@ 4.1 Experimental Design
We performed computational experiments to evaluate our approach to optimizing the Kolm-Pollak EDE against traditional facility location models in terms of both computational tractability and solution characteristics. Specifically, we compare the (KPL) model with both the p-median model (p-Med) and p-center model (p-Ctr). The p-median model represents the conventional approach focused on minimizing average travel distance, while the p-center model represents previous attempts to incorporate equity by protecting the worst-off residents.
To test these models at scale, we drew from two real-world transportation applications: optimizing access to grocery stores across the 500 largest U.S. cities (Horton et al., 2024), and locating early voting centers (polling locations) in five cities Skipper et al. (2025). The grocery store study provides extensive testing across diverse urban contexts and problem sizes, while the voting location study enables detailed analysis of solution properties and model behavior. For the grocery store application, we solved 4,500 instances total - 1,500 instances of each model - by optimizing the location of k = 1, 5, and 10 new facilities in each city while treating existing facilities as fixed. The polling application allowed all facilities to be relocated, providing a more challenging test of the models’ capabilities.
We implemented all models in Python using the optimization modeling language Pyomo (Hart et al., 2011; Bynum et al., 2021) and solved them using the mixed-integer linear programming solver Gurobi (Gurobi Optimization, LLC, 2023). For the grocery store instances, we used a high-performance computing cluster with an AMD 7502 CPU processor, allocating one of the 64 available cores and 512 GB memory to each instance. The New York City instances, which contained over 200 million binary variables, required additional memory and were solved using 2 TB RAM. The polling instances were solved using SCIP on a 2020 MacBook Pro with the Apple M1 chip and 16GB RAM to demonstrate performance on standard computing hardware and open-source solvers.
For the grocery store instances, we used Gurobi’s default optimality gap of 0.01%. For the polling instances, which involve relocating existing facilities and thus present a more challenging optimization problem, we set a time limit of three hours and an optimality gap of 0.5%. These parameters were chosen to balance solution quality with computational tractability based on preliminary testing.
4.2 Test Problems
4.2.1 Grocery Store Application
@@
Our primary test problem comes from a study analyzing food deserts across the 500 largest U.S. cities. That study examined strategic questions about improving equitable access to supermarkets, such as: given a limit of $k$ new stores, where should they be located to maximize equitable access? And what is the minimum number of new stores needed to achieve the U.S. average level of access? While those findings are presented in Horton et al. (2024), here we focus on the methodological aspects of solving these large-scale equitable facility location problems.
For each city, we incorporated:
@@
  • Census Block locations and populations from the 2020 U.S. Census (The Census Block is the highest spatial resolution that the U.S. Census reports data)
  • Existing supermarket locations from OpenStreetMap
@@
  • Potential new locations at Census Block Group centroids (A Block Group is the 2nd highest spatial resolution reported by the U.S. Census)
@@
  • Walking network distances calculated using the Open Source Routing Machine
This dataset generated substantial optimization problems. The largest case, New York City, involves 30,095 census blocks, 657 existing supermarkets, and 7,618 potential new locations, resulting in over 248 million binary variables. Los Angeles and Chicago each exceeded 100 million variables, while Houston and Philadelphia contained 61 and 28 million variables respectively (Table 9).
4.2.2 Polling Location Application
Our second application comes from our work with advocacy groups and election boards in preparation for the 2024 presidential election, where we developed and applied methods for optimizing early voting locations (Skipper et al., 2025). This study focused on questions such as evaluating the impact of historical changes to voting locations, identifying optimal placements for new locations, and analyzing access disparities between demographic groups.
These instances, summarized in Table 2, range from Salem’s 92 potential sites serving 3,102 census blocks to Baltimore’s 901 potential locations serving 9,238 blocks. The number of required facilities (k) varies from 8 to 256, reflecting each city’s actual polling requirements in 2016 or 2020.
Unlike the grocery store instances where existing facilities remained fixed, the polling location problems allowed all existing voting sites to potentially be relocated. This flexibility, combined with capacity constraints on the number of voters that could be assigned to each location, made these problems particularly challenging despite their relatively smaller size.
City Population Existing Supermarkets Number of Variables
New York, NY 8,784,592 657 248,975,936
Los Angeles, CA 3,849,235 537 108,306,340
Chicago, IL 2,733,239 316 106,859,845
Houston, TX 2,215,641 427 61,658,801
Philadelphia, PA 1,593,147 172 28,086,301
@@ 4.2.2 Polling Location Application
Table 9: Summary of data from five largest cities in the food deserts study.
4.3 Computational Results
4.3.1 Scalability Analysis
Figure 4 presents solve times for the food desert instances. The solve times correspond to the “wall time” for each instance, the length of time between when the solver is called and when a solution is returned. All instances solved to optimality using Gurobi’s default MIP gap of 0.01%. The top row compares solve times between (p-Med) and (KPL), while the second row compares (p-Ctr) and (KPL). The columns show results for k = 1, 5, and 10 new stores in each city.
Each point represents optimizing locations for a single city, with point size indicating the number of integer variables. The axes show solve times in seconds on a log scale, with (KPL) times on the horizontal axis and comparison model times on the vertical axis. Point color indicates which model solved faster.
We observe that (p-Med) and (KPL) achieved similar solve times across all instances, with each outperforming the other roughly half the time. This performance parity persisted even as the number of new stores increased. As expected, (KPL) typically solved significantly faster than (p-Ctr).
@@ 4.3.1 Scalability Analysis
The New York City instances, detailed in Table 10, provide specific insights into relative performance. While (KPL) took longer than (p-Med) in all cases, the gap was modest - 48% longer for k=1, but less than 0.05% difference for k=5 and k=10. Both models substantially outperformed (p-Ctr), which required approximately twice the solve time.
Figure 4: Food desert application solve times for (p-Med), (p-Ctr), and (KPL) on 500 largest U.S. cities for $k=1,5,10$ additional supermarkets. Note that solve times are displayed on a log scale; (p-Ctr) often takes roughly 10 times longer to solve than (KPL).
@@
4.3.2 Numerical Considerations
The (KPL) model has a larger range of objective coefficients than (p-Med), which can potentially cause numerical instability in optimization solvers. While this did not affect the food desert instances, Gurobi issued “large coefficient” warnings for polling Scenario 5. The coefficient range of $[2\times 10^{-6},~{}7\times 10^{28}]$ for (KPL)a expanded to $[2\times 10^{-6},~{}2\times 10^{42}]$ for (KPL)K, increasing solve time from 1.5 to 12.7 hours.
Two strategies can mitigate these numerical issues:
  • Reduce $|\epsilon|$ slightly - Section 3.3 showed optimal solutions are not highly sensitive to small changes in $\kappa$
  • Omit $y_{r,s}$ variables for assignments where $d_{r,s}$ exceeds some threshold $d_{max}$, eliminating clearly suboptimal possibilities
However, setting $d_{max}$ requires careful consideration to avoid excluding potentially beneficial assignments. The specific threshold choice depends on the problem context and should be validated through sensitivity analysis.
4.4 Properties of Optimal Solutions
@@
Each optimal solution to a food desert scenario defines a distribution of distances that residents must travel to reach their assigned supermarket. In this section, we compare statistics of the distributions of distances corresponding to optimal solutions to (KPL), (p-Med), and (p-Ctr).
@@
Figure 5 provides an overview of the mean distance traveled by residents to reach their closest grocery store in optimal solutions to (p-Ctr), (p-Med), and (KPL) with $\epsilon=-1$. The 500 largest cities are represented in each of the three charts, which display scenarios with $k\in\{1,5,10\}$ new stores. The cities are sorted along the horizontal axis according to the mean distance achieved by the optimal solution to (KPL), so the three colored shapes representing optimal solutions to (KPL), (p-Med), and (p-Ctr) for a single city appear in a vertical line. Notably, the average distance for (p-Ctr) is consistently higher than that of both (p-Med) and (KPL), while the average distances for (p-Med) and (KPL) closely align.
@@
Figure 5: Average walking distances for residents to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (KPL) models for 500 largest U.S. cities.
@@
Figure 6: Maximum walking distances for any resident to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (KPL) models for 500 largest U.S. cities.
@@
Figure 6 is similar to Figure 5 except that it presents the maximum (rather than mean) distance traveled in each optimal distribution of distances. The horizontal axis is sorted according to the maximum distance in an optimal solution to (KPL). (p-Med) consistently yields significantly higher maximum distances compared to both (p-Ctr) and (KPL). While (KPL) optimal solutions correspond to greater maximum distances when compared to (p-Ctr), (KPL) still results in a substantial improvement over the maximum distances in (p-Med) optimal solutions.
For the polling study, Table 11 provides four summary statistics (Kolm-Pollak EDE, mean, maximum, and standard deviation) for the distribution of distances corresponding to optimal solutions to the three models. The (p-Ctr) model applied to Scenario 5 did not solve to optimality within the three hour time limit. Figure 7 shows the distributions of distances the populations of each city must travel to reach their assigned polling location under optimal solutions to each of the models. In both the table and the figure, we see that the (KPL) solutions either very nearly match or improve on every summary statistic versus optimal solutions to the (p-Ctr) and (p-Med) models. The relatively poor performance of the (p-Ctr) solutions with respect to population access to amenities is apparent in Figure 7.
Figure 7: Cumulative percentage of population within $x$ meters of assigned polling location.
Scenario Model KP EDE Mean Max St. Dev.
(KPL) 6541.81 3087.90 18164.1 1734.13
1 (p-Med) 8074.17 2796.79 22780.1 1873.23
(p-Ctr) 11655.60 10281.16 18164.1 4249.91
(KPL) 2256.18 1573.41 6764.5 920.08
2 (p-Med) 2278.3 1569.93 6764.5 973.58
(p-Ctr) 5199.72 4805.95 6764.5 1576.32
(KPL) 2198.29 1605.90 6585.2 857.68
3 (p-Med) 2216.91 1601.98 6585.2 878.20
(p-Ctr) 4934.76 4549.96 6585.2 1497.15
(KPL) 5585.93 1073.50 11380.9 768.37
4 (p-Med) 6707.20 1072.08 14876.8 771.60
(p-Ctr) 5896.42 5311.81 11426.9 1994.96
(KPL) 571.21 454.21 2389.2 215.49
5 (p-Med) 2249.23 445.60 5218.0 241.31
(p-Ctr)
Table 11: Statistics of the driving distance distributions defined by the optimal solutions obtained by different models. Note that $\epsilon=-2$ both in the (KPL) model and in the calculation of the KP EDE. (KPL) provided a better solution with respect to maximum distance than (p-Ctr), but the difference is within the optimality gap of 0.5% that we set for these instances.
4.4.1 Average vs Equality
We have seen that (KPL) scales to very large practical problems, performing similarly to the (p-Med) model which is a common choice in facility location applications. However, one concern a city planner may have when choosing to use (KPL) over (p-Med) may be the fear of sacrificing average distance (overall welfare) in order to find an equitable solution (supporting those with the poorest access). Indeed, that is what we see in the (p-Ctr) solutions: the optimal (p-Ctr) travel distances are longer on average, often significantly longer. However, in the instances we have studied, (KPL) consistently offers the best of both worlds: near-optimal mean distances and a better outcome for the worst-off residents versus minimizing average distance directly.
This trend is highlighted In Table 12. On average across all cities, less than 10 meters in mean distance must be sacrificed in order to realize more than 400 meters of improvement in maximum distance. As expected, the improvements are more pronounced as more supermarkets are opened.
(KPL) versus (p-Med) optimal distributions $k=1$ $k=5$ $k=10$
Average increase in mean distance (meters) +8.859 +7.206 +5.7756
Average decrease in maximum distance (meters) -431.832 -489.088 -531.498
@@ 4.4.1 Average vs Equality
Table 12: Comparing decrease in maximum distance versus increase in mean distance for (KPL) versus (p-Med) optimal solutions. Averages are taken over the 500 “food deserts” cities when adding $k$ new supermarkets.

### v4

@@ 4.2 $\alpha$-Approximation Experiments: Grocery Stores
4.3 $\alpha$-Approximation Experiments: Election Polling Sites
For the grocery store instances, the accuracy of $\alpha_{1}^{in}$ decreases as the number of stores increases because the distribution of distances changes more as more stores locations are added. However, $\alpha_{2}^{in}$ remained a very good approximation, even as the number of stores increased. Moreover, (KPL)1 still resulted in high quality optimal solutions in almost all cases, in spite of the relatively poor performance of $\alpha_{1}^{in}$. To more extensively test our findings, we consider situations where the optimal distributions of distances is even further from the original distances.
For this analysis, we introduce another set of practical scenarios. In these problems, we seek to optimally relocate the same number of election polling sites that were used in either the 2016 or the 2020 presidential election in five large US cities. Data for the polling scenarios is summarized in Table 7. $|R|$ indicates the number of census blocks (residential areas), $|S|$ indicates the number of potential polling sites, and $k$ indicates the number of polling sites selected (which matches the number of polling sites used in the given election year). The potential polling sites have varying capacities so we use the (KPL) model with constraints (10). We set the desired aversion to inequality, denoted $\epsilon_{0}$, equal to 2-2 for all scenarios.
In the polling scenarios we expect the initial $\alpha_{1}^{in}$ to be a less-accurate approximation than in the grocery store instances because we allow all of the existing sites to move. This means that every resident may have a different distance in the new distribution. In contrast, with the grocery store examples, we do not remove any existing sites, so the travel distance for many residents will remain unchanged in the optimal distribution. We also expect the average distance, and the Kolm-Pollak EDE, to improve significantly from the original distribution to the optimal distribution, and this was indeed the case – we saw that the optimal Kolm-Pollak EDE represented an improvement by more than a mile in each of the five cities. Despite these potential concerns, we saw very similar behavior for the polling instances as we did for the grocery store instances.
In Table 8, we display all three $\alpha$ approximations and the “true” aversion to inequality values for (KPL)1 and (KPL)2. In Table 9, we present the raw and percent gaps between $\mathcal{K}_{1}^{*}$ and $\mathcal{K}_{2}^{*}$. In Table 8, we see that $\epsilon_{1}$ ranges between 1.56-1.56 and 0.55-0.55. The desired level of aversion (2-2) is not maintained by any of the optimal solutions. However, in spite of the differences between $\epsilon_{1}$ and the much more accurate $\epsilon_{2}$ (Table 8), the corresponding optimal Kolm-Pollak scores differ by no more than 1% in any of the five scenarios (Table 9).
Scenario City Election $|R|$ $|S|$ $k$ Number of Variables
1 Salem 2016 3102 92 8 285,476
2 Richmond 2016 6208 211 65 1,310,099
3 Atlanta 2016 8364 387 126 3,237,255
4 Cincinnati 2016 7670 517 256 3,965,907
5 Baltimore 2020 9238 901 201 8,324,339
Scenario $\epsilon_{0}$ $\epsilon_{1}$ $\epsilon_{2}$ $\alpha_{1}^{in}$ $\alpha_{1}^{out}=\alpha_{2}^{in}$ $\alpha_{2}^{out}$
1 -2 -1.56 -2.00 0.000192 0.000246 0.000246
2 -2 -0.90 -2.00 0.000212 0.000474 0.000474
3 -2 -0.74 -1.98 0.000179 0.000485 0.000488
4 -2 -0.55 -2.01 0.000171 0.000616 0.000614
5 -2 -1.42 -2.01 0.001277 0.001797 0.001789
Table 8: Exact and approximate values of $\alpha$ and the aversion to inequality, where $\epsilon_{0}$ is the “desired” aversion to inequality. The aversion to inequality associated with the second iteration of the model, $\epsilon_{2}$, is very accurate for these instances.
Scenario $\mathcal{K}_{1}^{*}$ $\mathcal{K}_{2}^{*}$ $\Delta$ gap
1 6541.81 6541.81 0.00 0.0000
2 2256.18 2259.49 3.31 0.0015
3 2198.29 2191.37 6.92 0.0032
4 5585.93 5571.38 14.54 0.0026
5 571.21 566.15 5.06 0.0089
Table 9: KP EDEs obtained via (KPL)1 are close to the best-know KP EDEs, which were obtained via (KPL)2. Distances in columns 2, 3, and 4 are measured in meters.
4.4 Summary
These experiments indicate that typically, optimal solutions are not highly sensitive to minor variations in $\alpha$, or equivalently, to varying levels of aversion to inequality. However, some practitioners may wish to have more control over the aversion to inequality represented by the optimal solution. To do this, our experiments suggest the following strategy: solve (KPL) and calculate the true aversion to inequality that the optimal solution represents. If the accuracy is not as desired, solve the model again using the distance distribution from the first solve to tune the $\alpha$ parameter.
@@ 5.1 Experimental Design
We performed computational experiments to evaluate our approach to optimizing the Kolm-Pollak EDE on large-scale data sets in terms of both computational tractability and solution characteristics. Specifically, we compare the (KPL) model with the p-centdian model (p-Ctdn), the p-median model (p-Med), and the p-center model (p-Ctr). We chose these models as our benchmarks for the following reasons:
  • (p-Ctdn) represents a measure that explicitly balances efficiency with equity and still retains hope of scaling to the size instances we wish to solve. While there are many other metrics that account for this balance, it is clear they are not appropriate for even our moderate-sized instances (recall the detailed discussion in Section 2).
  • While only accounting for efficiency and not equity, (p-Med) is known to scale well to large models. We use this as a benchmark on the solve time of our models.
  • Furthermore, (p-Med) and (p-Ctr) may be viewed as the extreme cases of both (p-Ctdn) and (KPL), and a goal of both models is to appropriately balance the two. We consider optimal solutions to these extremes so that we may evaluate how well (p-Ctdn) and (KPL) balance the corresponding efficiency and fairness measures.
5.1.1 Data
@@
To test these models at scale, we drew from a real-world transportation application: optimizing access to grocery stores across the 500 largest U.S. cities. In Horton et al. (2024), we use this data to answer strategic questions for specific cities, such as: given a limit of $k$ new stores, where should they be located to maximize equitable access? And what is the minimum number of new stores needed to achieve the U.S. average level of access? In contrast, here, we demonstrate the methodological aspects of solving these large-scale equitable facility location problems. In doing so, we are able to extensively test our model on diverse urban contexts and problem sizes. Moreover, the only aspect of our data that is specific to supermarket access, as opposed to locating any general amenity, is the locations we chose to fix open.
For each city, we incorporate:
@@
  • Census Block locations and populations from the 2020 U.S. Census (the Census Block is the highest spatial resolution that the U.S. Census reports data);
  • existing supermarket locations from OpenStreetMap (OpenStreetMap contributors, 2017);
@@
  • potential new locations at Census Block Group centroids (a Block Group is the 2nd highest spatial resolution reported by the U.S. Census);
@@
  • walking network distances calculated using the Open Source Routing Machine (Luxen and Vetter, 2011).
5.1.2 Computing Environments
To fully explore the scalability and solution quality of our model we performed two distinct sets of experiments. In the first, we used a 2019 Dell XPS 13 9380 laptop with the Intel Core i7-8565U-8 chip and 16GB RAM to demonstrate performance on standard computing hardware. Here, our goal was to compare computational time and the maximum size of solvable instances when using (KPL) with that achievable by the p-centdian approach.
We optimized the locations of $k\in\{1,5,10\}$ new facilities in each of the 500 cities while treating existing facilities as fixed open. We sought to solve all instances on four models: the (KPL) model with $\epsilon=-1$, and (p-Ctdn) with $\gamma\in\{0.1,0.5,0.9\}$. We chose to consider only one value of $\epsilon$ as preliminary experiments showed very little variation in the solve time for values of $\epsilon$ within the commonly-accepted interval [2,0.5][-2,-0.5].
Our dataset included some very large optimization models, with the largest corresponding to New York City, which included 30,094 census blocks and 8,274 destinations, resulting in over 249 million binary variables in the (KPL) models. In Table 10 we summarize the problem data for five cities of interest. The first three cities correspond to the largest problem instances in the data set, as indicated by the first column which reports the rank of each city (among all 500 cities) according to the number of variables in the (KPL) models for that city. At rank 61 with around 1.5 million variables, Glendale, AZ is the highest ranked city for which all three (KPL) models (for $k\in\{1,5,10\}$) could be solved on the laptop. At rank 136 with around 0.5 million variables, Garden Grove, CA is the highest ranked city for which all nine (p-Ctdn) models (for $\gamma\in\{0.1,0.5,0.9\}$ and $k\in\{1,5,10\}$) could be solved on the laptop.
To further test the (KPL) model without the memory restrictions of a standard laptop, we ran all 500 cities (1,500 instances of the model) with $\epsilon=-1$ on a high-performance computing cluster with an AMD 7502 CPU processor, allocating one of the 64 available cores and 512 GB memory to each instance. The New York City instances required additional memory and were solved using 2 TB RAM. In this setting, we were able to solve (KPL) on every instance for all 500 cities to optimality. We also solved all instances on both the (p-Med) and (p-Ctr) models using the cluster.
In all cases, we implemented the models in Python using the optimization modeling language Pyomo (Hart et al., 2011; Bynum et al., 2021) and solved them using the mixed-integer linear programming solver Gurobi (Gurobi Optimization, LLC, 2025), using the default optimality gap of 0.01%.
Rank City Population —R— —S— Number of Variables
1 New York, NY 8,784,592 30,094 8,274 249,006,030
2 Los Angeles, CA 3,849,235 23,038 4,701 108,306,340
3 Chicago, IL 2,733,239 33,011 3,237 106,859,845
61 Glendale, AZ 241,320 2,249 706 1,588,500
136 Garden Grove, CA 171,455 951 608 578,816
@@ 61 Glendale, AZ 241,320 2,249 706 1,588,500
Table 10: Summary of data from the three largest cities in the food deserts study, plus the largest instances solved on a standard laptop for (KPL) and (p-Ctdn) respectively.
5.2 Computational Results: Scalability Analysis
5.2.1 Comparison with p-centdian Using a Personal Laptop
The personal laptop solved all nine (p-Ctdn) models for 348 cities and all three (KPL) models for 428 cities. There were an additional six cities where (KPL) was solved for $k=1$, but not for $k=5$, $k=10$, or both. The laptop encountered memory errors when attempting to load models corresponding to the other 66 cities. There were also many instances (86 cities) for which the laptop was able to load the data and the (p-Ctdn) model, but the kernel crashed during presolve due to memory issues.
Figure 7 presents performance profiles showing the solve times for solved instances. The horizontal axis is on a logarithmic scale and displays the solve time (in seconds); this corresponds to the “wall time” for each instance, i.e., the length of time between when the solver is called and when a solution is returned. The vertical axis shows the number of cities that were solved to optimality within the given number of seconds. We observe that (KPL) out-performs (p-Ctdn) by around a factor of 10 for all values of $k$ and $\gamma$.
Figure 7: Performance profiles showing food desert application solve times on a laptop for (p-Ctdn) and (KPL) on 428 U.S. cities for $k\in\{1,5,10\}$ additional supermarkets. (The parameter is $\gamma$ and refers to the ($p$-Ctdn) instances). The shorter lines for the (p-Ctdn) model indicate that fewer instances of that model solved on the laptop, and the position of the lines indicate that for the problems that (p-Ctdn) succeeded at solving, it was around 10 times slower than (KPL).
5.2.2 Comparison with p-median and p-center Using a High-Performance Computing Cluster
To overcome the memory limitations on a standard laptop, we solved (KPL) with $\epsilon=-1$ on the high-performance computing cluster for all 500 of the largest US cities. Here, we compare solve time with (p-Med), i.e., (p-Ctdn) with $\gamma=0$. This represents the conventional approach focused on efficiency, which we know scales well to large instances. At the other end of the spectrum, we compare solve time with (p-Ctr), i.e., (p-Ctdn) with $\gamma=1$.
@@ 5.2.2 Comparison with p-median and p-center Using a High-Performance Computing Cluster
Observing the performance profiles in Figure 8, we see that (p-Med) and (KPL) achieved similar solve times across all instances, and this performance parity persisted as the number of new stores increased. As expected, (KPL) typically solved significantly faster than (p-Ctr). The New York City instances, detailed in Table 11, provide specific insights into relative performance. While (KPL) took longer than (p-Med) in all cases, the gap was modest, especially as the number of additional stores increased – less than 0.05% difference for $k=5$ and $k=10$. Both models substantially outperformed (p-Ctr), which required approximately twice the solve time.
Figure 8: Performance profiles showing food desert application solve times on the High Performance Computing Cluster for (KPL), (p-Med) and (p-Ctr) on all 500 U.S. cities for $k=1,5,10$ additional supermarkets.
@@
5.3 Computational Results: Properties of Optimal Solutions
@@
Each optimal solution to a problem instance defines a distribution of distances that residents must travel to reach their assigned (in this case, nearest) supermarket. To visually compare the solution quality of (KPL) with that of (p-Ctdn), we plot the maximum value and the mean value of the distributions given by the respective optimal solutions. In particular, we wish to compare the mean value of each optimal distribution with the best possible mean value, obtained by solving (p-Med), and similarly, we compare the maximum value of each optimal distribution with the optimal maximum value, obtained by solving (p-Ctr).
@@
Figure 9(a) provides an overview of how the mean distance traveled by residents to reach their closest grocery store in an optimal solution of (KPL) compares with optimal solutions to (p-Ctr) and (p-Med). The 500 cities are represented on the horizontal axis in each of the three charts, which display scenarios with $k\in\{1,5,10\}$ new stores. They are sorted according to the mean distance achieved by the optimal solution to (KPL), so the three colored shapes representing optimal solutions to (KPL), (p-Med), and (p-Ctr) for a single city appear in a vertical line. Figure 9(b) is similar, but compares (p-Ctdn) to (p-Ctr) and (p-Med) for $\gamma\in\{0.1,0.5,0.9\}$ for the 348 cities for which we have (p-Ctdn) results.
Notably, the average distance for (p-Ctr) is consistently much higher than that of both (p-Med) and (KPL), while the average distances for (p-Med) and (KPL) closely align. This general trend holds for (p-Ctdn) with $\gamma=0.1$, although interestingly, it seems that an even smaller value of $\gamma$ would be necessary to align with (p-Med) to the same extent as (KPL).
@@
(a) Average walking distances for residents to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (KPL) models for 500 largest U.S. cities.
@@
(b) Average walking distances for residents to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (p-Ctdn) models for 350 largest U.S. cities.
Figure 9:
@@
Figures 10(a) and 10(b) are similar to Figures 9(a) and 9(b) except that they present the maximum (rather than mean) distance traveled in each optimal distribution of distances. The horizontal axis is sorted according to the maximum distance in an optimal solution to (KPL) (resp. (p-Ctdn)).
As expected, we observe that (p-Med) consistently yields significantly higher maximum distances compared to (p-Ctr), (KPL) and p-(Ctdn). While (KPL) optimal solutions correspond to greater maximum distances when compared to (p-Ctr), they still result in a substantial improvement over the maximum distances in (p-Med) optimal solutions. For $\gamma=0.1$, (p-Ctdn) demonstrates a similar balance, while still remaining closer to the (p-Ctr) solutions. For larger values of $\gamma$, (p-Ctdn) aligns extremely closely with (p-Ctr).
(a) Maximum walking distances for any resident to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (KPL) models for 499 largest U.S. cities. Note: one outlier with a large (p-Med) maximum value was removed from the $k=1$ store chart to improve readability.
(b) Maximum walking distances for any resident to nearest supermarket in optimal solutions of (p-Med), (p-Ctr), and (p-Ctdn) models for 350 U.S. cities.
Figure 10:
We wrap up this section with a small table of statistics on the set of cities that were solved for both (p-Ctdn) and (KPL). In particular, we demonstrate the tradeoff between improving the maximum distance traveled and increasing the average travel distance versus the optimal (p-Med) solution. We can see that even with $\gamma=0.1$, which heavily weights improving the average distance versus the maximum distance, the (p-Ctdn) model is more successful at decreasing the maximum distance, while the (KPL) is more successful at maintaining a near-optimal average experience while addressing equity across the distribution. Of course, the (p-Ctdn) could be tuned to focus more on average distance by choosing a smaller value for $\gamma$.
$k=1$ $k=5$ $k=10$
(KPL) versus (p-Med)
Average increase in mean distance (meters) +9.63 +5.26 +3.08
Average decrease in maximum distance (meters) -319.26 -376.03 -429.45
(p-Ctdn) versus (p-Med)
Average increase in mean distance (meters) +26.81 +33.17 +26.80
Average decrease in maximum distance (meters) -722.84 -1187.79 -1226.48
@@
Table 12: Comparing decrease in maximum distance versus increase in mean distance for (KPL) with $\epsilon=-1$ and (p-Ctdn) with $\gamma=0.1$ relative to (p-Med) optimal solutions. Averages are taken over the 348 food desert cities for which the (p-Ctdn) models solved successfully.
5.3.1 Differences between (KPL) and (p-Ctdn)
We have seen that optimal solutions to both (KPL) and (p-Ctdn) can be interpreted as navigating the trade-off between efficiency and fairness by balancing the mean (p-Med) and maximum (p-Ctr) objectives. However, it is important to note that while both metrics can be interpreted in this way, the Kolm-Pollak EDE is able to distinguish between distributions even when p-centdian is not.
Viewed like this, we see that (KPL) may in fact serve as a better proxy to the general Discrete Ordered Median Problem (DOMP) than (p-Ctdn) when the goal is to solve larger instances than can be handled using the general framework. Nuanced solutions that address equity across the entire distribution, such as those attained by specially constructed versions of the DOMP, cannot be achieved by (p-Ctdn) because the measure only takes into account the maximum and mean values of a distribution. Hence, in the context of the DOMP, all distances that are not the maximum are forced to have equal weight. This means that in practice, the corresponding facility location problem may often have multiple optimal solutions – two distributions with equal mean and maximum values cannot be differentiated in (p-Ctdn). In contrast, the general DOMP is able to appropriately distinguish between distributions with equal mean and maximum values by assigning suitable weights to every distance in the distribution. Example 5 demonstrates that the Kolm-Pollak EDE similarly distinguishes between such distributions in a reasonable way. Moreover, as we have seen, (KPL) retains the desirable computational properties that allow us to solve instances much larger than if we were to use a true DOMP.
Example 5.
Consider a simple scenario with three residential areas, each with a population of one. We are interested in the value of the Kolm-Pollak EDE and the p-centdian metric for a family of distance distributions, the details of which are provided in Table 13.
Residential Area Population Distance to Assigned Facility (as a function of $\Delta$)
1 1 $0+\Delta$
2 1 $80-\Delta$
3 1 100100
Table 13: We consider a family of distance distributions, parameterized by $\Delta\in[0,80]$, for three residential areas. Note that the maximum of each distribution is equal to 100 and does not depend on $\Delta$, moreover, the mean of each distribution is equal to $180/3=60$, and also does not depend on $\Delta$.
The family is constructed such that the mean value and the maximum value do not depend on the parameter $\Delta$. Hence, the value of the p-centdian remains constant as we vary $\Delta$ for any fixed choice of $\gamma$. However, while the mean and maximum values do not change, the family is constructed so that the equality in the distribution does vary with $\Delta$. With $\Delta=0$, or $\Delta=80$, the distribution has maximum inequality, conversely, when $\Delta=40$, the distribution is significantly more fair.
Figure 11 displays the value of the p-centdian for $\gamma=0.1,0.25,0.5$ and the Kolm-Pollak EDE for $\epsilon=-0.5,-1,-2$ as $\Delta\in[0,80]$ varies. As expected, p-centdian does not distinguish between the quality of the distributions and remains constant, whereas the Kolm-Pollak EDE is able to reflect our preference for distributing the 80 units of distance between residential areas 1 and 2 as equally as possible.
Figure 11: Comparing the Kolm-Pollak and p-centdian metrics for the family of distributions as $\Delta$ varies from 0 to 80. The Kolm-Pollak (blue, dashed) is always able to account for the difference in equality as $\Delta$ varies, whereas the p-centdian (red, dotted) is not, remaining constant for all choices of $\gamma$.